Queen, Inc., has a total debt ratio of .42.

a. What is its debt-equity ratio? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.)
b. What is its equity multiplier? (Do not round intermediate calculations. Round your answer to 2 decimal places, e.g., 32.16.)
Answer #1

Total Debt Ratio = Total Debt/Total Assets = 0.42

a) Debt Equity Ratio = 0.42/(1-0.42) = 42/58 = 0.72

b) Equity Multiplier = Total Assets/Equity = 1/)1-0.42) = 1/0.58 = 1.72

Best of Luck. God Bless
Please Rate Well